### v2

@@ 4 Generalization to moments selection
Theorem 2 shows how our framework generalizes to two related problems (i) choosing the weighting matrix $W$ for estimation and (ii) moment selection. The key insight is to show that the regret is quasi-convex under (approximate) linearity in the moments.
5 Extensions
5.1 Local asymptotics in the presence of non-linearities
This subsection extends our framework for (i) nonlinear moment functions $g(\theta)$ used in GMM in Section 4 and (ii) a smooth, potentially nonlinear estimand $\tau(\theta)$.
@@ 5.1 Local asymptotics in the presence of non-linearities
Assume the sample moments obey root–$n$ scaling, i.e., $\mathbb{V}(\bar{g}_{\Sigma})=O(n^{-1})$. Let $b=b_{n}$ be a sequence where retain the same misspecification structure as in Section 4: the moment vector has mean $b=b_{n}$ with $(b_{n})_{j}=0$ for $j\in\mathcal{I}$ (experiment-based moments) and $(b_{n})_{\mathcal{I}^{c}}$ potentially nonzero. Our key assumption is a local condition on the moments bias of the form
@@
This condition builds on the local asymptotic frameworks in Andrews et al. (2020) and Armstrong and Kolesár (2021). In their context, $b_{n}||^{2}=1/n$ matching the rate of convergence of the variance. In our case, $b_{n}$ can grow faster, slower or at the same rate of the standard error, encompassing $b_{n}||^{2}=1/n$ as a special case.
@@
$\sqrt{n}\big(\hat{\theta}(W,\Sigma)-\theta\big)=-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,\sqrt{n}\big(\bar{g}_{\Sigma}-b_{n}\big)+o_{p}(1),$
so on the original (unscaled) scale the mean shift is $(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,b_{n}$ and the remainder is $o_{p}(n^{-1/2})$. The condition $\|b_{n}\|^{2}\sqrt{n}\to 0$ ensures that higher order terms are negligible under sufficient smoothness conditions, justifying the use of the linear expansion (5) for our design analysis. Because $b_{n}\to 0$, the Jacobian $\Lambda$ can be consistently estimated by evaluating the derivative of $g(\theta)$ at a preliminary observational estimator $\tilde{\theta}^{\mathrm{obs}}\rightarrow_{p}\theta$, assuming such estimates are available from an observational study (Andrews et al., 2020).
@@
$\tau(\hat{\theta})-\tau(\theta)=\omega(\theta)^{\top}(\hat{\theta}-\theta)\;+\;\tfrac{1}{2}\,(\hat{\theta}-\theta)^{\top}H_{\tau}(\bar{\theta})\,(\hat{\theta}-\theta),$
for some $\bar{\theta}$ between $\hat{\theta}$ and $\theta$, and $H_{\tau}$ is the Hessian. It follows that we can write
@@
Remark 2 (Rescaling and change of variables linking Example 3).
@@
An equivalent formulation uses $\sqrt{n}$–rescaled coordinates (as in Example 3). Define the rescaled moment vector $\tilde{g}:=\sqrt{n}\,\bar{g}$ and the rescaled estimation error $\tilde{\theta}\equiv\sqrt{n}\,(\hat{\theta}-\theta)$. Then
$\mathbb{V}(\tilde{g})=\Sigma=O(1),\qquad\mathbb{E}[\tilde{g}]=b_{n}\ \ \text{with}\ \ b_{n}\equiv\sqrt{n}\,b,$
so $b_{n}$ is the asymptotic bias and $\Sigma$ the asymptotic variance. The local misspecification condition in (9) is equivalent to $\frac{\|b_{n}\|^{2}}{\sqrt{n}}\;\to\;0,$ which coincides with Example 3. ∎
5.2 Confidence interval length as a regret criterion
In this subsection we consider an alternative loss based on the length of confidence intervals (CIs), focusing on Setting 1 for simplicity. For a candidate design $(\mathcal{E},\Sigma(\mathcal{E}))$, shrinkage weights $\gamma$, and a bias vector $b\in\mathbb{R}^{p}$, define the two–sided, $(1-\eta)$ lower and upper bounds for $\tau(\theta)$ by
$\ell_{b}(\mathcal{E},\Sigma,\gamma)$ $\equiv\hat{\tau}_{\gamma}\;-\;(\omega\circ(1-\gamma))^{\top}b\;-\;z_{1-\eta/2}\,\sqrt{\alpha(\mathcal{E},\Sigma,\gamma)},$
$u_{b}(\mathcal{E},\Sigma,\gamma)$ $\equiv\hat{\tau}_{\gamma}\;-\;(\omega\circ(1-\gamma))^{\top}b\;+\;z_{1-\eta/2}\,\sqrt{\alpha(\mathcal{E},\Sigma,\gamma)},$
where $z_{1-\eta/2}$ is the standard normal $(1-\eta/2)$ quantile.
An audience, indexed by a worst–case bias radius $B\geq 0$, forms the worst–case CI
$L_{B}(\mathcal{E},\Sigma,\gamma)~\equiv~\left[\inf_{\|b\|_{\infty}\leq B}\ell_{b}(\mathcal{E},\Sigma,\gamma),\;\sup_{\|b\|_{\infty}\leq B}u_{b}(\mathcal{E},\Sigma,\gamma)\right],$
where we write $|L_{B}(\mathcal{E},\Sigma,\gamma)|$ for its total length. The proportional regret of $(\mathcal{E},\Sigma,\gamma)$ is
$\tilde{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)~\equiv~\sup_{B\geq 0}\;\frac{|L_{B}(\mathcal{E},\Sigma,\gamma)|}{\inf_{(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})\in\mathcal{D}}|L_{B}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})|}\!.$
Theorem 3.
Consider Setting 1 and let Assumption 2 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)$,
[…]
Proof.
See Appendix B.1
Theorem 3 shows that the objective function is the same as the MSE-objective function up to a monotonic (square-root) transformation which does not affect the minimizer.
5.3 Vector valued estimands
Next, we extend our framework to vector value estimands of the form $\tau(\theta)\in\mathbb{R}^{q}$ for $q\geq 1$.
Setting 3 (Vector-valued estimands).
Consider a vector-valued target estimand with entries $\tau^{\ell}(\theta)\equiv(\omega^{\ell})^{\top}\theta$ for $\ell=1,\dots q$, and $\omega^{\ell},\theta\in\mathbb{R}^{k}$.
Given $(\mathcal{E},\Sigma)$, and observational and experimental estimates $\tilde{\theta}^{\mathrm{obs}},\tilde{\theta}^{\mathrm{exp}}$, consider a class of linear plug-in estimators to allow the optimal shrinkage weights to depend on the specific entry of $\tau$.
$\hat{\tau}^{\ell}_{\gamma}\equiv\tau^{\ell}(\hat{\theta}^{\ell}(\gamma)),\qquad\hat{\theta}^{\ell}_{j}(\gamma)=\begin{cases}\gamma_{j}^{\ell}\,\tilde{\theta}_{j}^{\mathrm{exp}}+(1-\gamma_{j}^{\ell})\,\tilde{\theta}_{j}^{\mathrm{obs}},&j\in\mathcal{E},\\[2.0pt] \tilde{\theta}_{j}^{\mathrm{obs}},&j\notin\mathcal{E},\end{cases}$
where $\gamma=(\gamma^{\ell}_{j})_{\ell\in\{1,\cdots,q\},j\in\mathcal{E}}\in\mathbb{R}^{q\times|\mathcal{E}|}$ are shrinkage weights that can be an (implicit) function of $(\mathcal{E},\Sigma)$. Let $\hat{\theta}(\gamma)\equiv(\hat{\theta}^{1}(\gamma),\dots,\hat{\theta}^{q}(\gamma))\in\mathbb{R}^{p\times q}$ and $\hat{\tau}_{\gamma}\equiv(\hat{\tau}^{1}_{\gamma}(\hat{\theta}^{1}(\gamma)),\dots,\hat{\tau}^{q}_{\gamma}(\hat{\theta}^{q}(\gamma)))^{\top}\in\mathbb{R}^{q}$.
Setting 3 is a flexible generalization of Setting 1, allowing each entry of $\tau$ to have its own vector of shrinkage weights.
Assumption 4 (First-order estimation error).
For any admissible $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$, assume for all $\ell=1,\dots,q$
$\tau^{\ell}(\theta)-\tau^{\ell}\!\big(\hat{\theta}^{\ell}(\gamma)\big)=(\omega^{\ell})^{\top}\,\big(\theta-\hat{\theta}^{\ell}(\gamma)\big),$ (10)
for a known weighting vector $\omega^{\ell}\in\mathbb{R}^{p}$, with $|\omega^{\ell}_{i}|\in(0,\infty)$, for all $i,\ell$.
We begin by generalizing the mean squared error to higher-dimensional parameters. For a given design $(\mathcal{E},\Sigma)$ and shrinkage weights $\gamma$, define the MSE at a fixed observational-bias vector $b$ as
$\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)\;=\;\mathbb{E}_{\mathcal{E},\Sigma,b}\!\left[\|\widehat{\tau}_{\gamma}-\tau\|_{2}^{2}\right],$ (11)
where $\mathbb{E}_{\mathcal{E},\Sigma,b}$ denotes expectation under the data-generating process implied by $(\mathcal{E},\Sigma(\mathcal{E}))$ and observational bias $b$, and $\|\cdot\|_{2}$ denotes the $L^{2}$ Euclidean norm.
We introduce more compact notation before proceeding to the main result. For a given $(\mathcal{E},\gamma)$ define
$\tilde{\omega}^{\ell}_{\text{obs},j}(\mathcal{E},\gamma)=\begin{cases}\omega^{\ell}_{j},&j\notin\mathcal{E},\\ \omega^{\ell}_{j}(1-\gamma^{\ell}_{j}),&j\in\mathcal{E},\end{cases}\qquad\tilde{\omega}^{\ell}_{\text{exp},j}(\mathcal{E},\gamma)=\begin{cases}0,&j\notin\mathcal{E},\\ \omega^{\ell}_{j}\gamma^{\ell}_{j},&j\in\mathcal{E},\end{cases}$
and let $\tilde{\omega}^{\ell}(\mathcal{E},\gamma)\equiv\begin{pmatrix}\tilde{\omega}^{\ell}_{\text{obs}}(\mathcal{E},\gamma)\\[2.0pt] \tilde{\omega}^{\ell}_{\text{exp}}(\mathcal{E},\gamma)\end{pmatrix}\in\mathbb{R}^{\,p+|\mathcal{E}|}.$ We can write $\hat{\tau}^{\ell}_{\gamma}=\tilde{\omega}^{\ell}(\mathcal{E},\gamma)^{\top}\,\hat{\theta}$, where $\hat{\theta}=(\tilde{\theta}^{\text{obs}\top},\tilde{\theta}^{\text{exp}\top})^{\top}\in\mathbb{R}^{p+|\mathcal{E}|}$. Stacking across the $\ell$ index gives the estimator for the full vector $\tau$: $\hat{\tau}_{\gamma}=\tilde{\omega}(\mathcal{E},\gamma)\,\hat{\theta}$, where the $\ell^{th}$ row of $\tilde{\omega}(\mathcal{E},\gamma)\in\mathbb{R}^{q\times(p+|\mathcal{E}|)}$ is $\tilde{\omega}^{\ell}(\mathcal{E},\gamma)^{\top}$.
The estimator’s variance and worst–case squared bias divided by $B^{2}$ are
$\check{\alpha}(\mathcal{E},\Sigma,\gamma)$ $\equiv\text{Trace}\left(\tilde{\omega}(\mathcal{E},\gamma)\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)^{\top}\right),$ (12)
$\check{\beta}(\mathcal{E},\gamma)$ $\equiv\sum_{\ell=1}^{L}\Big(\,(\omega^{\ell})^{\top}\bar{v}(\gamma)-(\omega^{\ell}_{\,\mathcal{E}})^{\top}\bar{v}(\gamma)_{\,\mathcal{E}}+\,(1-\gamma^{\ell})^{\top}(\bar{v}(\gamma)_{\,\mathcal{E}}\circ\omega_{\,\mathcal{E}})\,\Big)^{\!2},$
where $\omega^{\ell}_{\mathcal{E}}$ is the subvector of $\omega^{\ell}$ corresponding to the entries indexed by the experiment $\mathcal{E}$, and
$\bar{v}(\gamma)\in\operatorname*{arg\,max}_{\{v\in\mathbb{R}^{p}:|v_{j}|\leq 1\forall j\}}\sum_{\ell=1}^{L}\Big(\,(\omega^{\ell})^{\top}v-(\omega^{\ell}_{\,\mathcal{E}})^{\top}v_{\,\mathcal{E}}+\,(1-\gamma^{\ell})^{\top}(v_{\,\mathcal{E}}\circ\omega_{\,\mathcal{E}})\,\Big)^{\!2},$
where the dependence of $\bar{v}(\gamma)$ on $\mathcal{E}$ is omitted for conciseness and $\circ$ denotes the component-wise multiplication. The smallest worst-case variance and bias are defined as
$\check{\alpha}^{\star}=\min_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\alpha(\mathcal{E},\Sigma,\gamma),\quad\check{\beta}^{\star}=\min_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\beta(\mathcal{E},\gamma).$
As before, we are interesting in minimizing
$\check{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)\equiv\sup_{B\geq 0}\frac{\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}{\inf_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}.$
The result below shows how our results directly extend to this case.
Theorem 4.
Consider Setting 3 and let Assumption 4 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)$,
$\check{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)\;=\;\max\!\left\{\frac{\check{\alpha}(\mathcal{E},\Sigma,\gamma)}{\check{\alpha}^{\star}},\ \ \frac{\check{\beta}(\mathcal{E},\gamma)}{\check{\beta}^{\star}}\right\}.$
Proof.
See Appendix B.2

### v3

@@ 4 General framework for moment selection
Theorem 2 shows how our framework generalizes to (i) choosing the moments and the estimator (via $W$); and (ii) arbitrary norms characterizing the ambiguity set. The key insight is that the regret is quasi-convex under (approximate) linearity in the moments relative to the dual norm of the ambiguity set.
This generalization encompasses, in addition to choosing which experiment to run, problems that involve whether to collect additional data (e.g., through surveys or data acquisition) since these can be directly incorporated as part of the moment conditions.
Example 7 (Weighted bias).
Let $\tau(\theta)\in\mathbb{R}$ and consider an ambiguity set of the form $\tilde{\mathcal{B}}(B)\;\equiv\;\Big\{b\in\mathbb{R}^{p}:|b_{j}|\leq k_{j}B\;\;\text{for all }j=1,\dots,p\Big\},$ for known rescaling factors $k$. Then the corresponding bias regret is $\tilde{\beta}/\tilde{\beta}^{\star}$ with $\tilde{\beta}(\mathcal{E},\gamma)\;\equiv\;(\sum_{j\notin\mathcal{E}}k_{j}|\omega_{j}|\;+\;\sum_{j\in\mathcal{E}}k_{j}|\omega_{j}|\big|1-\gamma_{j}\big|)^{2},\quad\tilde{\beta}^{*}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}(\sum_{j\notin\mathcal{E}}k_{j}|\omega_{j}|)^{2},$ and the overall regret equals $\max\!\left\{\,\frac{\alpha(W,\Sigma)}{\alpha^{\star}}\,,\ \frac{\tilde{\beta}(W)}{\tilde{\beta}^{\star}}\,\right\}$. ∎
5 Confidence sets, incomplete models and extensions
5.1 Confidence sets and partially identified $\tau$
Next, we generalize our framework to minimizing confidence interval length and partially identified $\tau(\theta)$. For simplicity, we let $\tau(\theta)\in\mathbb{R}$. Researchers may not know $\tau(\cdot)$ exactly and only know functions $\bar{\tau},\underline{\tau}$ with
$\bar{\tau}(\theta)=\bar{\omega}^{\top}\theta,\qquad\underline{\tau}(\theta)=\underline{\omega}^{\top}\theta,\quad\tau(\theta)\in[\underline{\tau}(\theta),\bar{\tau}(\theta)]$
where linearity can be justified as a first-order linear approximation for $\bar{\tau}(\theta)$ and $\underline{\tau}(\theta)$ under the local asymptotics in Section 5.3 for smooth functions $\bar{\tau},\underline{\tau}$. [^12] When the envelope for $\tau(\theta)$ is non-smooth, we recommend replacing it with smooth upper and lower envelopes. In this case, following Section 5.3, $\bar{\omega}$ and $\underline{\omega}$ can be interpreted as the gradients of these smooth bounds evaluated at the observational estimates. We implicitly assume that at least one of $\bar{\omega}$ and $\underline{\omega}$ differ from the zero vector to avoid trivial solutions.
For a generic weight vector $\omega$ let $\alpha_{\omega}(W,\Sigma),\beta_{l,\omega}(W)$ as in Definition 3.
Confidence intervals under bias.
For a candidate design $(W,\Sigma)$ and a bias vector $b\in\mathbb{R}^{p}$, define two-sided $(1-\eta)$ lower and upper bounds for $\tau(\theta)$ by
$\ell_{b}(W,\Sigma)$ $\equiv\underline{\tau}(\hat{\theta}(W))\;-\;\underline{\omega}^{\top}\Gamma_{\Lambda}(W)b\;-\;z_{1-\eta/2}\,\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)},$
$u_{b}(W,\Sigma)$ $\equiv\bar{\tau}(\hat{\theta}(W))\;-\;\bar{\omega}^{\top}\Gamma_{\Lambda}(W)b\;+\;z_{1-\eta/2}\,\sqrt{\alpha_{\bar{\omega}}(W,\Sigma)},$
where $z_{1-\eta/2}$ is the standard normal $(1-\eta/2)$ quantile and $\hat{\theta}(W)$ is the GMM estimator in (17). These bounds adjust for the bias $\omega^{\top}\Gamma_{\Lambda}(W)b$ in $\omega^{\top}\hat{\theta}(W)$ for each envelope separately.
As before, the Jacobian $\Lambda$ is known; when estimated see Section 5.3 and Remark 5.
An audience, indexed by a worst-case bias radius $B\geq 0$, forms a worst-case confidence interval (i.e., a bias-aware confidence interval a la Armstrong and Kolesár (2018)) by choosing the most conservative endpoints over $b$ in the ambiguity set $\mathcal{B}_{l}(B)$ in (16):
$L_{l,B}(W,\Sigma)~\equiv~\left[\inf_{b\in\mathcal{B}_{l}(B)}\ell_{b}(W,\Sigma),\;\sup_{b\in\mathcal{B}_{l}(B)}u_{b}(W,\Sigma)\right],$
and we denote by $|L_{l,B}(W,\Sigma)|$ its total length.
The researcher does not need to know the audience’s specific choice of $B$ and can simply report the point estimates and associated standard errors; the audience (indexed by $B$) can then form their own bias-aware confidence set from these statistics. The researcher would like to minimize the expected length of the confidence interval formed by the audience.
An oracle that knows $B$ minimizes the worst-case expected length of the audience’s confidence length relative to the size of the sharp identified set (worst-case under the audience’s prior $b\in\mathcal{B}_{l}(B)$). Its (ex-ante) expected loss is defined as
$\mathcal{L}_{l,B}(W,\Sigma)~\equiv~\sup_{b_{0}\in\mathcal{B}_{l}(B)}\Big\{\mathbb{E}_{W,\Sigma,b_{0}}\big[|L_{l,B}(W,\Sigma)|\big]-\big(\bar{\tau}(\theta)-\underline{\tau}(\theta)\big)\Big\}.$ (21)
Here $\mathcal{L}_{l,B}(W,\Sigma)$ measures how informative the expected confidence interval is, relative to the length of the identified set $\bar{\tau}(\theta)-\underline{\tau}(\theta)$, worst-case over $b_{0}$ in $\mathcal{B}_{l}(B)$. Here, $b_{0}$ determines the expectation of $\hat{\theta}(W)$, and ultimately the length of the bias-aware identified set.
As in previous sections, we define the corresponding regret as
$\tilde{\mathcal{R}}_{l}(W,\Sigma)~\equiv~\sup_{B\geq 0}\;\frac{\mathcal{L}_{l,B}(W,\Sigma)}{\inf_{(W^{\prime},\Sigma^{\prime})\in\mathcal{D}^{\prime}}\mathcal{L}_{l,B}(W^{\prime},\Sigma^{\prime})}\!.$
Example 8 (Confidence set in a complete model).
Consider a complete model with $\bar{\tau}(\theta)=\underline{\tau}(\theta)=\tau(\theta)$ and $\bar{\omega}=\underline{\omega}=\omega$. Then we can show
$\mathcal{L}_{l,B}(W,\Sigma)=2\,z_{1-\eta/2}\sqrt{\alpha_{\omega}(W,\Sigma)}+2\,B\,\sqrt{\beta_{l,\omega}(W)}.$
In particular, when the parameter is point identified, the excess confidence interval length depends only on the variance index $\alpha_{\omega}$ and the bias index $\beta_{l,\omega}$, corresponding to standard bias-aware confidence intervals (Armstrong and Kolesár, 2018).
Example 9 (Simple incomplete model).
[…]
t=0t=0t=1t=1t=2t=2t=3t=3 Design
Choose $(W,\Sigma)$
Reporting
Compute $\hat{\theta}(W)$
[…]
Report $\alpha_{\underline{\omega}}$, $\alpha_{\bar{\omega}}$
Audience
Observe $(W,\Sigma)$
Choose $B\in\mathbb{R}$
Form $L_{l,B}(W,\Sigma)$
Ex-post loss
Excess length
$|L_{l,B}(W,\Sigma)|-(\bar{\tau}(\theta)-\underline{\tau}(\theta))$
Figure 2: Timeline of design, reporting, audience confidence set, and loss.
Regret optimal design (and estimator).
Define
$A(W,\Sigma)\equiv z_{1-\eta/2}\big(\sqrt{\alpha_{\bar{\omega}}(W,\Sigma)}+\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)}\big),\quad C_{l}(W)\equiv\sqrt{\beta_{l,\bar{\omega}}(W)}+\sqrt{\beta_{l,\underline{\omega}}(W)}+\sqrt{\beta_{l,\bar{\omega}-\underline{\omega}}(W)},$
and the corresponding oracle solutions
$A^{\star}\equiv\inf_{(W,\Sigma)\in\mathcal{D}^{\prime}}A(W,\Sigma),\qquad C_{l}^{\star}\equiv\inf_{(W,\Sigma)\in\mathcal{D}^{\prime}}C_{l}(W).$
The first component $A$ depends on the variance of the estimated parameters (and envelopes) and the second component $C_{l}$ depends on the bias. We now show how our results directly extend to this scenario.
Theorem 3 (Regret for confidence interval length).
Let Assumptions 2 and 3 hold and consider an estimator as in (17). Then, for any $(W,\Sigma)\in\mathcal{D}^{\prime}$,
$\tilde{\mathcal{R}}_{l}(W,\Sigma)=\max\Big\{\frac{A(W,\Sigma)}{A^{\star}},\;\frac{C_{l}(W)}{C_{l}^{\star}}\Big\}.$
Proof.
See Appendix A.2. ∎
Theorem 3 shows that the confidence-interval-length regret can be decomposed into a variance component $A(W,\Sigma)$ and a bias component $C_{l}(W)$, each divided by its smallest feasible value. This result mimics Theorem 2 allowing for partially identified models. It provides a ready-to-use expression for experimental design (and moment selection).
The following corollary shows that whenever the model is point identified the solutions that minimize the regret relative to confidence interval length coincides with the MSE optimal solution we derived in Section 4.
Corollary 1 (Equivalence of MSE and confidence-interval length optimal solutions).
Let Assumptions 2 and 3 hold. Let $\alpha,\beta_{l},\alpha^{\star},\beta_{l}^{\star}$ be as in Definition 3, and consider an estimator as in (17). Suppose $\bar{\omega}=\underline{\omega}=\omega$ (complete model). Then, for any $(W,\Sigma)$,
$\arg\min_{(W,\Sigma)\in\mathcal{D}^{\prime}}\mathcal{R}_{l}(W,\Sigma)~=~\arg\min_{(W,\Sigma)\in\mathcal{D}^{\prime}}\tilde{\mathcal{R}}_{l}(W,\Sigma),$
where $\mathcal{R}_{l}(W,\Sigma)$ denotes the MSE-based regret in Equation (19).
Proof.
See Appendix A.3. ∎
In summary, when $\tau(\theta)$ is point identified, minimizing the regret of the confidence interval length leads to the same optimal design as minimizing MSE-based regret. In partially identified models, Theorem 3 shows that the confidence interval length maintains the same bias-variance decomposition, a desirable property of the MSE solution in Section 4.
5.2 Optimization with partial knowledge on the bias bound
In this section we extend our results to settings where researchers have prior information about the bias upper bound. Following the framework in Section 4, suppose researchers know that $B\leq\bar{B}$ for a known upper bound $\bar{B}$, and consider an arbitrary (possibly weighted) norm $\|\cdot\|_{l}$ on the bias. The theorem below establishes the resulting characterization of the regret.
Theorem 4 (Regret characterization).
Consider the setting in Section 4. Let Assumptions 2 and 3 hold, and let $\mathcal{R}$ be as defined in Equation (19), with $\alpha_{\Omega},\beta_{l,\Omega},\alpha_{\Omega}^{\star},\beta_{l,\Omega}^{\star}$ as in Definition 3. Consider an estimator as in Equation (17). Then
$\sup_{B\in[0,\bar{B}]}\ \frac{\sup_{b\in\mathcal{B}_{l}(B)}\mathrm{MSE}_{b}(W,\Sigma)}{\inf_{(W^{\prime},\Sigma^{\prime})\in\mathcal{D}^{\prime}}\sup_{b^{\prime}\in\mathcal{B}_{l}(B)}\ \mathrm{MSE}_{b^{\prime}}(W^{\prime},\Sigma^{\prime})}\;=\;\max\!\left\{\,\frac{\alpha_{\Omega}(W,\Sigma)}{\alpha_{\Omega}^{\star}}\,,\ \frac{\alpha_{\Omega}(W,\Sigma)+\bar{B}^{2}\beta_{l,\Omega}(W)}{\min_{W^{\prime},\Sigma^{\prime}}\big\{\alpha_{\Omega}(W^{\prime},\Sigma^{\prime})+\bar{B}^{2}\beta_{l,\Omega}(W^{\prime})\big\}}\,\right\}.$ (22)
Proof.
See Appendix A.1. ∎
Theorem 4 shows that when an upper bound $\bar{B}$ on the bias is known, the worst-case proportional regret over all $B\in[0,\bar{B}]$ is again the maximum of two terms. The first term, $\alpha_{\Omega}(W,\Sigma)/\alpha_{\Omega}^{\star}$, is the familiar variance regret. The second term compares the MSE evaluated at the largest feasible bias level, $\bar{B}$, to the smallest achievable MSE at the same bias level across all feasible designs and weight matrices. The optimization problem is therefore analogous to our baseline case, but with the “bias” component of regret evaluated at $B=\bar{B}$ rather than in the limit as $B\to\infty$.
From an interpretation standpoint, this extension interpolates between two polar cases. When $\bar{B}$ is small, the second term is close to the variance ratio, and designs that minimize variance dominate. As $\bar{B}$ increases, the second term places more weight on robustness to misspecification, and the objective approaches the fully robust case studied in our baseline results. In practice, the theorem implies that any prior information that narrows $B$ to $0,\bar{B}]$ can be incorporated simply by replacing the worst-case bias component in our original criterion with the MSE at $B=\bar{B}$, while keeping the same quasi-convex structure.
5.3 Local asymptotics with non linear target
Next, we extend our framework for (i) nonlinear moment functions $g(\theta)$ used in Section 4 and (ii) a smooth, potentially nonlinear estimand $\tau(\theta)$. Although we focus on known $\tau(\cdot)$, similar reasoning applies to partially identified sets in Section 5.1, omitted for brevity.
@@ 5.3 Local asymptotics with non linear target
Assume the sample moments obey root-$n$ scaling, i.e., $\mathbb{V}(\bar{g}_{\Sigma})=O(n^{-1})$. Let $b=b_{n}$ be a sequence where we retain the same misspecification structure as in Section 4: the moment vector has mean $b=b_{n}$ with $(b_{n})_{j}=0$ for $j\in\mathcal{I}$ (experiment-based moments) and $(b_{n})_{\mathcal{I}^{c}}$ potentially nonzero. Our key assumption is a local condition on the moments bias of the form
@@
This condition builds on the local asymptotic frameworks in Andrews et al. (2020) and Armstrong and Kolesár (2021), where $b_{n}||^{2}=1/n$ matching the rate of convergence of the variance. In our case, $b_{n}$ can grow faster, slower or at the same rate of the standard error (i.e., $b_{n}\propto n^{-\alpha},\alpha>1/4$), encompassing $b_{n}||^{2}=1/n$ as a special case.
@@
$\sqrt{n}\big(\hat{\theta}(W,\Sigma)-\theta\big)=-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,\sqrt{n}\big(\bar{g}_{\Sigma}-b_{n}\big)+o_{p}(1).$
The condition $\|b_{n}\|^{2}\sqrt{n}\to 0$ ensures that higher order terms are negligible under sufficient smoothness conditions, justifying the use of the linear expansion (17) for our design analysis. Because $b_{n}\to 0$, the Jacobian $\Lambda$ can be consistently estimated by evaluating the derivative of $g(\theta)$ at a preliminary observational estimate $\tilde{\theta}^{\mathrm{obs}}\rightarrow_{p}\theta$ (Andrews et al., 2020).
@@
Remark 5 (Ex-post valid coverage with estimated $\Lambda$).
When some components of $\hat{\theta}$ are based on observational data and $\Lambda$ is estimated from the same observational sample used to select the design, the selection of the experiment and the construction of $\hat{\theta}$ both depend on the same data. Unless the observational sample is (asymptotically) much larger than the experimental sample, this case can lead to selection-induced distortions in the distribution of $\hat{\theta}$ and invalidate standard Gaussian confidence intervals (see, e.g., Andrews et al., 2024). A simple way to avoid this problem is to split the observational data into two independent subsamples (independent of the experimental samples): one used to estimate $\Lambda$ and choose the experiment at the design stage, and a second used, together with the experimental data to estimate the observational and experimental components of $\hat{\theta}$. Under this sample-splitting scheme the experimental and observational data used for inference are independent of the design choice, so standard CLT-based confidence intervals retain their nominal coverage. ∎
Remark 6 (Equivalent rescaling).
@@
An equivalent formulation uses $\sqrt{n}$–rescaled coordinates (as in Example 4). Define the rescaled moment vector $\tilde{g}:=\sqrt{n}\,\bar{g}$ and the rescaled estimation error $\tilde{\theta}\equiv\sqrt{n}\,(\hat{\theta}-\theta)$. Then $\mathbb{V}(\tilde{g})=\tilde{\Sigma}=O(1),\mathbb{E}[\tilde{g}]=\tilde{b}_{n}\ \ \text{with}\ \ \tilde{b}_{n}\equiv\sqrt{n}\,b,$ so $\tilde{b}_{n}$ is the asymptotic bias and $\tilde{\Sigma}$ the asymptotic variance. Local misspecification in (23) is equivalent to $\frac{\|\tilde{b}_{n}\|^{2}}{\sqrt{n}}\;\to\;0,$ which coincides with notation in Example 4. ∎